Falcon Corporation purchased a depreciable asset for $840,000 on January 1, 2015. The estimated salvage value is $84,000, and the estimated total useful life is 9 years. The straight-line method is used for depreciation. In 2018, Falcon changed its estimates to a total useful life of 5 years with a salvage value of $140,000. What is 2018 depreciation expense?              
Answer #1

Annual depreciation = (Cost price – Residual value)/Useful life

= (840,000 - 84,000)/9

= $84,000

Accumulated depreciation from 2015 to 2017 = 84,000 x 3

= $252,000

Book value of asset on January 1, 2018 = Cost price - Accumulated depreciation from 2015 to 2017

= 840,000 - 252,000

= $588,000

New depreciable cost = Book value of asset – Revised residual value

= 588,000 - 140,000

= $448,000

Revised annual depreciation = New depreciable cost/Remaining useful life

= 448,000/2

= $224,000

Thus, 2018 depreciation expense = $224,000

Kindly comment if you need further assistance. Thanks